OpenStudy (anonymous):

PLEASE SOLVE AND SHOW STEPS!!!!! (x-4)-5(x-4)^.5=6

OpenStudy (anonymous):

\[(x-4)-5(x-4)^{.5}=6\] Isolate \[(x-4)^{.5}\] on the left. Square both sides. You should get ... my answer's ugly, though, are you sure you've written it out correctly. Or have I made a mistake?

OpenStudy (anonymous):

gimmick is to write \[(x-4)-5\sqrt{x-4}-6=0\] and then since \[(x-4)=\sqrt{x-4}^2\] you have a quadratic equation in \[\sqrt{x-4}\]

OpenStudy (anonymous):

i.e. put \[z=\sqrt{x-4}\]and then solve \[z^2-5x-6=0\] \[(z-6)(z+1)=0\] \[z=6\] so \[\sqrt{x-4}=6\] \[x-4=36\] \[x=40\]

OpenStudy (anonymous):

ohhhhh... duh

OpenStudy (anonymous):

notice you can ignore \[\sqrt{x-4}=-1\] because it is not possible. another solution would be to write \[-5\sqrt{x-4}=6-x+4=10-x\] and then square both sides to get \[25(x-4)=100-20x+x^2\] \[25x-100=100-20x+x^2\] \[x^2-45x+200=0\] \[(x-40)(x-5)=0\] \[x=40\] or \[x=5\] but the check will show that 5 does not work